OpenStudy (anonymous):

HELP!! suppose that G(x) = log (base 4) (2x+2)-2 a) what is the domain of G? b) what is g(31)? what point is on the graph of G? c) If G(x)=3, what is x? what point is on the graph of G? d) what is the zero of G?

OpenStudy (anonymous):

\[G(x)=\log_4(2x+2)-2\] domain -> all values of "x" when G can be calculated..

OpenStudy (anonymous):

so, when do you think you cannot calculate "G"? hint: can you calculate log of "0" or any thing negative?

OpenStudy (anonymous):

can not calculate anything that's below zero right??

OpenStudy (anonymous):

"log" of "0" and negative numbers is not defined

OpenStudy (anonymous):

so, we write, \[2x+2>0\]

OpenStudy (anonymous):

simplify that and what is the inequality for "x"?

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

you write it as x > -1

OpenStudy (anonymous):

yeah just noticed

OpenStudy (anonymous):

great. so, the domain for G(x) is, any "x" greater than 1 i.e., from "1" ( but not equal to 1) to infinity

OpenStudy (anonymous):

so, domain = \((1,\infty)\)

OpenStudy (anonymous):

okay!

OpenStudy (anonymous):

how would you find b? since i cant do log base 4 on calculator

OpenStudy (anonymous):

yes remember the change of base rule? \[\log_a(x)=\frac{\log_{10}x}{\log_{10}a}\]

OpenStudy (anonymous):

a=4 log (base10) you have it on your calcultor

OpenStudy (anonymous):

we didn't learn that yet.

OpenStudy (anonymous):

hmm. that is the way you can calculate logarithm to any base.

OpenStudy (anonymous):

so it would be log base 10 31/ log base 10 4?

OpenStudy (anonymous):

oh no wait.. my bad I took the problem wrong.. and you didnot even correct me!!! \[G(x)=\log_4(2x+x)-2\] this is what you have, right?

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok, so doing the steps for (a) again \[2x+x>0\]

OpenStudy (anonymous):

are you sure there are no exponents in ther?

OpenStudy (anonymous):

no exponents

OpenStudy (anonymous):

ok.. so, proceeding the same way now, what is your domain?

OpenStudy (anonymous):

what happens to the -2?

OpenStudy (anonymous):

-2 is not in the paranthesis of the "log" so, it just shifts the function G(x) up,down along the Y-axis

OpenStudy (anonymous):

wait im sorry it is actually right the first time I did it wrong on here! it is actually (2x+2)-2. I am SUPER SORRY!!!!!!!!

OpenStudy (anonymous):

write it exactly using the equation editor (button is below this reply box)

OpenStudy (anonymous):

\[G(x)=\log _{4}(2x+2)-2\]

OpenStudy (anonymous):

so, by co-incidence, I got it right!!!

OpenStudy (anonymous):

yes! sorry i missed typed it on here!

OpenStudy (anonymous):

so would it be the same for b?

OpenStudy (anonymous):

yes. would not change. replace "x" by 31 and evaluate

OpenStudy (anonymous):

I got 2.48

OpenStudy (anonymous):

does that look correct?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

2*31+2=?

OpenStudy (anonymous):

44?

OpenStudy (anonymous):

\(2\times31+2=?\)

OpenStudy (anonymous):

64. so the base rule dosent work since I used it and got the 2.48

OpenStudy (anonymous):

wait now do \[\log_4(64)\]

OpenStudy (anonymous):

using base-change rule.. it always works as long as you change to valid bases

OpenStudy (anonymous):

oh okay so now it would be \[\left(\begin{matrix}\log _{10}64 \\ \log _{10}4\end{matrix}\right)\]

OpenStudy (anonymous):

perfect.. (if you mean division)

OpenStudy (anonymous):

hahaha yes

OpenStudy (anonymous):

so the answer would be three!

OpenStudy (anonymous):

makes more sense now

OpenStudy (anonymous):

great, now finally, G(31)=3-2

OpenStudy (anonymous):

will be 1

OpenStudy (anonymous):

bingo yay..

OpenStudy (anonymous):

so i am super confused on how to do the last 2 problems

OpenStudy (anonymous):

how would you start c?

OpenStudy (anonymous):

wait.. with (b).. what is the point?

OpenStudy (anonymous):

you put it as (x-value,y-value) so the point is (31,1)

OpenStudy (anonymous):

point is (31, 1)

OpenStudy (anonymous):

yes. for G(3) put x=3 and evaluate.. show me your steps so I know what you are doing

OpenStudy (anonymous):

so it would be \[3=\log _{4}(2x+2)-2\]

OpenStudy (anonymous):

close... you still have the "x".. rewrite the given function as it is with all "x" replaced by "3"

OpenStudy (anonymous):

including the "G" part

OpenStudy (anonymous):

I thought that it was g(x)=3

OpenStudy (anonymous):

nonono .. just the "x"

OpenStudy (anonymous):

\[G(x=3)=\log_4[2(3)+2]-2\]

OpenStudy (anonymous):

but we are trying to find for x

OpenStudy (anonymous):

ooh right.. my bad you are perfectly going well.. sorry to break your flow

OpenStudy (anonymous):

lmao its okay i was kinda stuck anyways. do you add 2 to 3 to get \[5=\log _{4}(2x-2)\]

OpenStudy (anonymous):

good.. now, how do you get rid of the "log (base 4)"?

OpenStudy (anonymous):

do you do ln of 4

OpenStudy (anonymous):

opposite of log is exponentiation and vice versa so, you exponent the equation to base "4" since that is the base of your "log" \[4^5=(2x+2)\]

OpenStudy (anonymous):

okay so then do you -2 and then divide by 2

OpenStudy (anonymous):

perfect

OpenStudy (anonymous):

511?

OpenStudy (anonymous):

yes!!

OpenStudy (anonymous):

the point would be (511,3)

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

so how would we find the zero

OpenStudy (anonymous):

the zero of G

OpenStudy (anonymous):

it means, for what value of "x" is G(x)=0?

OpenStudy (anonymous):

so, set G(x)=0 and solve for "x" like part (c)

OpenStudy (anonymous):

i got -31/32

OpenStudy (anonymous):

nope.. steps?

OpenStudy (anonymous):

i did 0=log base 4 (2x+2)-2 2=log base 4 (2x+2) 4^2=2x+2 7=x opps did it wrong the first time. so 7 would be the answer

OpenStudy (anonymous):

right.

OpenStudy (anonymous):

haha thank you super much! is there chance you can help me with 2 more problems?

OpenStudy (anonymous):

I have a good friend who is very nice.

OpenStudy (anonymous):

post your question as a new one and I'll refer her there